OpenStudy (anonymous):

HELP Find the area of the region that lies inside both curves r = 8sin(2theta), r = 8sin(theta)

OpenStudy (roadjester):

let me guess, third year calculus?

OpenStudy (anonymous):

yeah why?

OpenStudy (roadjester):

It was torture for me :) let me see if I can help, I need some scratch paper though

OpenStudy (anonymous):

ok thanks man its so tough

OpenStudy (roadjester):

i can relate, gimmie a sec

OpenStudy (anonymous):

ok

OpenStudy (roadjester):

have you done double integrals?

OpenStudy (anonymous):

Not yet

OpenStudy (roadjester):

dang, okay gotta think of how to do this using single...

OpenStudy (anonymous):

we haven't gotten that far we are in chapter 9 with areas of regions and in curves

OpenStudy (roadjester):

give me as much info as you can then

OpenStudy (anonymous):

well the question is just right there lol, idk what else you want

OpenStudy (anonymous):

we have done polar coordinates, taylor series , maclaurin series

OpenStudy (roadjester):

okay good, you can't do this in rectangular, polar is your best, no only shot unless you want roots. it's also cleaner. I just need to remember how to do a single integral sadly. I haven't done Calc 3 in a long time

OpenStudy (anonymous):

Yeah I hear you, I'm looking at this problem and I don't even know where to start, if you can show me step by step what you do, that would be so helpful

OpenStudy (roadjester):

\[8\sin(2\theta)=8\sin(\theta)\] \[\sin(2\theta)=\sin(\theta)\] \[2\sin(\theta)\cos(\theta)=\sin(\theta)\] \[2\sin(\theta)\cos(\theta)-\sin(\theta)=0\] \[\sin(\theta)(2\cos(\theta)-1)=0\] \[\sin(\theta)=0\] and \[2\cos(\theta)-1=0\] Solve for theta and these are your limits of integration.

OpenStudy (roadjester):

Notice how I cancelled the 8 but didn't cancel the sin(theta). You never want to cancel out anything that isn't a constant since it might be a potential solution.

OpenStudy (anonymous):

ok I see

OpenStudy (roadjester):

Can you pick it up from there since you've done polar coordinates? I don't know if you've done integration using polar yet or not.

OpenStudy (anonymous):

so its 2 cos(theta) = 1, so cos(theta) = 1/2?

OpenStudy (roadjester):

But believe me, you'll need a strong grip on polar since double and triple integrals build on that. Cylindrical is like a second layer to polar and spherical...well just be strong in polar and double and triple will be easier for you

OpenStudy (roadjester):

Yes

OpenStudy (anonymous):

OK so i have sin(theta) = 0 and cos(theta) = 1/2

OpenStudy (roadjester):

yep, and notice how both functions are even( you're in polar not rectangular) so you need to take into account the negative answer

OpenStudy (anonymous):

Ok so now what?

OpenStudy (roadjester):

what are your values for theta?

OpenStudy (anonymous):

0 and 1/2?

OpenStudy (roadjester):

come on, your values for theta, not sin(theta) and cos(theta). Think trig or geometry. This is calculus so if you're willing to take calc, especially calc 3, you should know how to take the inverse of a trig function.

OpenStudy (anonymous):

I thought you meant those values though

OpenStudy (roadjester):

theta refers to the angle. What are the angles? Stop and don't think calculus right now. Go back to your geometry or trig. How did you find an angle using trig functions like sine and cosine?

OpenStudy (anonymous):

The circle, so wouldn't it be 30 degrees

OpenStudy (roadjester):